good morning. I need help in knowing the amount of mortage they obtained, may you please help I don't understand...................this is my question......
Dan and Dawn purchased a house for $69,500. They had to make a down
payment of 20%. How much was the amount of the mortgage they obtained?

Answer by LinnW(1048)

The down payment + mortgage = $69,500
Since the down payment is 20%,
this means the down payment is 20% of $69500
= 0.20*69500 = 13900 = $13900
Since the down payment + mortgage = 69500
13900 + mortgage = 69500
subtract 13900 from each side
mortgage = 69500 - 13900 = 55600 = $55,600
and
13900 + 55600 = 69500